OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±7).

OpenStudy (mertsj):

This hyperbola opens up and down as opposed to right and left.

OpenStudy (mertsj):

The center is half way between the vertices so the center is (0,0)

OpenStudy (mertsj):

a=2, c=7 a^2+b^2=c^2

OpenStudy (mertsj):

\[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (anonymous):

hank you so much :)

OpenStudy (mertsj):

yw